OpenStudy (anonymous):

An object having a mass of 2.0 kilograms falls from a height of 15 meters. What is its kinetic energy when it hits the ground?

OpenStudy (anonymous):

KE = PE KE = 1/2 mv^2 PE = mgh PE = 2kg * 9.81 m/s^2 * 15m = 294 Joules Since potential energy is the same as kinetic energy or can interchange with it (converts back and forth) KE = 294 J

OpenStudy (theeric):

Which is to say that: The kinetic energy that it \(\sf will\ have\) is the potential energy. Basically, you say that the ground is \(0\) potential, meaning the energy you can potentially have is \(0\) (because it's not going to fall if it's already on the ground). Theennn, you look at where it is, and see how much energy it can potentially have (if it gets to the ground). The formula for this is three variable. \({\rm{Potential\ Energy} }=m\ g\ h\equiv\rm mass\times gravitational\ acceleration\times height\)